# Question-Answering Dense Video Events

### v4

@@ 1. Introduction
Multimodal Large Language Models (MLLMs) (Alayrac et al., 2022; Li et al., 2023; Maaz et al., 2024; Zhang et al., 2023b; Lin et al., 2023; Google, 2024) are highly capable at question-answering (QA) for single-event videos (Xu et al., 2017; Ye et al., 2017; Jang et al., 2017). Such videos are short (3 $\sim$ 20 seconds) and the QA factor single, global events, e.g. “who did what”. Yet, real-world videos are long and feature a complex overlay of dense events. Consider the 2-minute video of a motorcycle activity shown in Figure 1 (Top). A variety of questions can be asked about this video, each pertaining to an individual event but involving different participants and durations interspersed throughout the video.
@@ 3.1. Dataset Construction
For each question, we first obtain 10 candidate answers that are semantically reasonable to answer the question. We achieve this by retrieving the top 10 most similar questions about the same video and borrowing their answers to form the candidate answer list. If there are fewer than 10 QAs of the same video, we fetch the most similar QAs from other videos. The question similarity is calculated by the cosine score of their CLIP embeddings. Then, we pick two candidate answers that are mostly relevant to the video but are not present within the question time span as two distractor answers. This is achieved by using the segment-irrelevant video content (frames outside the question time span) to retrieve its top-two closest candidate answers (again via cosine similarity of their CLIP embeddings). Similarly, we pick one candidate answer that is mostly close to the video content specified by the question time span but is not the same as the correct answer. Finally, we add one more randomly sampled candidate answer from the whole answer set to ensure that it does not overlap the existing distractor answers and the correct answer (via simple lexical proximity).
Manual QA Checking and Curation. As all QAs are automatically generated, we further conduct manual check and correction to the test QAs. Our criteria are to ensure that the distractor answers: 1) are not potentially correct; 2) can logically respond to the given question; 3) are distinct from one another; 4) are closely related to the video content. The checking is done by 35 annotators over the course of 267 hours. Around 74% of QA pairs are modified, with examples of the final QAs shown in Figure 1 and 3.
@@ 4.3. Temporal Event Memory
To capture long-term dependencies, we design a memory module to contextualize the event captions while storing the visual and event representations. The LLM is are asked to refine the captions by prompting “… given a set of event captions {E} and a question {q} of a video, you are required to refine each caption by incorporating contextual information from all the other captions and question via analyzing the overall narratives, identifying relevant context and incorporate context with coherence…”. We also curate examples to perform in-context learning for the LLMs before the actual generation. We additionally prompt GPT-4o to articulate all the events into a synopsis $e_{y}$ to serve as a global event for the entire video. Our specific prompts are given in Table 4. Consequently, we obtain $E^{\prime}=\{E_{s}^{{}^{\prime}},E_{m}^{{}^{\prime}},E_{l}^{{}^{\prime}},e_{y}\}$, in which the events at each level are enhanced with long-range temporal dependency.
@@ 5.3. Ablation Studies
Table 9 presents the ablation results of the three major components in DeVi on DeVE-QA. Substituting the hierarchical dense captioning with a strategy of naively captioning every frame (as in LLoVi (Zhang et al., 2024b)) decreases QA and GQA accuracy by 4.9% and 4.4% respectively. Further in-depth analyses in Table 10 verify that hierarchical dense captioning facilities answering questions about denser event videos. Without it, the QA performance drops by 5.7% on the video subset with denser events, whereas the performance only drops by 2.3% and 3.7% on the video subsets with sparse events (1 to 2 events). This further demonstrates its ability to capture specific information from multiple and complicated events.
@@
Qualitative analyses of other baseline models also highlight their limitations. For instance, even though VideoLLaMA2 (Cheng et al., 2024) detected the roughly correct time span of the target event (Figure 7(a) row 3), its ineffective processing of temporal dynamics leads to its false understanding of the key event content and finally derive a wrong answer. LongVA (Zhang et al., 2024c) is good at handling long videos, yet it falls short in coping with obfuscating content with similar characteristics of objects/persons and recognizing the correct event span for other confusing events (Figure 7(b) row 3).

### v5

@@ 1. Introduction
Multimodal Large Language Models (MLLMs) (Alayrac et al., 2022; Li et al., 2023; Maaz et al., 2024; Zhang et al., 2023b; Lin et al., 2023; Google, 2024) are highly capable at question-answering (QA) for single-event videos (Xu et al., 2017; Ye et al., 2017; Jang et al., 2017). Such videos are short (3 $\sim$ 20 seconds) and the QAs factor single, global events, e.g. “who did what”. Yet, real-world videos are long and feature a complex overlay of dense events. Consider the 2-minute video of a motorcycle activity shown in Figure 1 (Top). A variety of questions can be asked about this video, each pertaining to an individual event but involving different participants and durations interspersed throughout the video.
@@ 3.1. Dataset Construction
For each question, we first obtain 10 candidate answers that are semantically reasonable to answer the question. We achieve this by retrieving the top 10 most similar questions about the same video and borrowing their answers to form the candidate answer list. If there are fewer than 10 QAs of the same video, we fetch the most similar QAs from other videos. The question similarity is calculated by the cosine score of their CLIP embeddings. Then, we pick two candidate answers that are mostly relevant to the video but are not present within the question time span as two distractor answers. This is achieved by using the segment-irrelevant video content (frames outside the question time span) to retrieve its top-two closest candidate answers (again via cosine similarity of their CLIP embeddings). Similarly, we pick one candidate answer that is mostly close to the video content specified by the question time span, but is not the same as the correct answer. Finally, we add one more randomly sampled candidate answer from the whole answer set to ensure that it does not overlap with the existing distractor answers and the correct answer (via simple lexical proximity).
Manual QA Checking and Curation. As all QAs are automatically generated, we further conduct a manual check and correction to the test QAs. Our criteria are to ensure that the distractor answers: 1) are not potentially correct; 2) can logically respond to the given question; 3) are distinct from one another; 4) are closely related to the video content. The checking is done by 35 annotators over the course of 267 hours. Around 74% of QA pairs are modified, with examples of the final QAs shown in Figure 1 and 3.
@@ 4.3. Temporal Event Memory
To capture long-term dependencies, we design a memory module to contextualize the event captions while storing the visual and event representations. The LLM is asked to refine the captions by prompting “… given a set of event captions {E} and a question {q} of a video, you are required to refine each caption by incorporating contextual information from all the other captions and question via analyzing the overall narratives, identifying relevant context and incorporate context with coherence…”. We also curate examples to perform in-context learning for the LLMs before the actual generation. We additionally prompt GPT-4o to articulate all the events into a synopsis $e_{y}$ to serve as a global event for the entire video. Our specific prompts are given in Table 4. Consequently, we obtain $E^{\prime}=\{E_{s}^{{}^{\prime}},E_{m}^{{}^{\prime}},E_{l}^{{}^{\prime}},e_{y}\}$, in which the events at each level are enhanced with long-range temporal dependency.
@@ 5.3. Ablation Studies
Table 9 presents the ablation results of the three major components in DeVi on DeVE-QA. Substituting the hierarchical dense captioning with a strategy of naively captioning every frame (as in LLoVi (Zhang et al., 2024b)) decreases QA and GQA accuracy by 4.9% and 4.4%, respectively. Further in-depth analyses in Table 10 verify that hierarchical dense captioning facilities answering questions about denser event videos. Without it, the QA performance drops by 5.7% on the video subset with denser events, whereas the performance only drops by 2.3% and 3.7% on the video subsets with sparse events (1 to 2 events). This further demonstrates its ability to capture specific information from multiple and complicated events.
@@
Qualitative analyses of other baseline models also highlight their limitations. For instance, even though VideoLLaMA2 (Cheng et al., 2024) detected the roughly correct time span of the target event (Figure 7(a) row 3), its ineffective processing of temporal dynamics leads to its false understanding of the key event content and finally results in a wrong answer. LongVA (Zhang et al., 2024c) is good at handling long videos, yet it falls short in coping with obfuscating content with similar characteristics of objects/persons and recognizing the correct event span for other confusing events (Figure 7(b) row 3).